# CombineHarvesterFlow: Joint Probe Analysis Made Easy with Normalizing Flows

### v1

@@ Abstract
We show how to efficiently sample the joint posterior of two non-covariant experiments with a large set of nuisance parameters. Specifically, we train an ensemble of normalizing flows to learn the the posterior distribution of both experiments. Once trained, we can use the flows to draw $\mathcal{O}(10^{9})$ samples from the joint posterior of two independent cosmological measurements in seconds – saving up to $\mathcal{O}(1)$ ton of $\text{CO}_{2}$ per Monte Carlo run. Using this new technique we find joint constraints between the Dark Energy Survey $3\times 2$ point measurement, South Pole Telescope and Planck CMB lensing and a BOSS direct fit full shape analyses, for the first time. We find $\Omega_{\rm m}=0.32^{+0.01}_{-0.01}$ and $S_{8}=0.79^{+0.01}_{-0.01}$. We release a public package called CombineHarvesterFlow [^1] https://github.com/pltaylor16/CombineHarvesterFlow which performs these calculations.
@@ I Introduction
Analyses of data from the Atacama Cosmology Telescope (ACT) Qu et al. (2024), the Baryon Oscillation Spectroscopic Survey (BOSS) Alam et al. (2021), the Dark Energy Survey (DES) Abbott et al. (2022), the Dark Energy Spectroscopic Instrument (DESI) Adame et al. (2024), the extended Baryon Oscillation Spectroscopic Survey (eBOSS)) Alam et al. (2021), Hyper Suprime-Cam (HSC) More et al. (2023), the Kilo-degree Survey (KiDS) Heymans et al. (2021), Planck Aghanim et al. (2020) and the South Pole Telescope (SPT) Bianchini et al. (2020), have provided some of the tightest cosmological constraints. Future measurements with Euclid, the Nancy Grace Roman Space Telescope, the Rubin Observatory, the Simons Observatory and CMB-S4 will tighten these constraints even further.
Meanwhile there has been a proliferation of cosmological likelihood packages to perform inference with these datasets (e.g. MontePython Brinckmann and Lesgourgues (2019), CosmoMC Lewis (2013), COSMOLIKE Krause and Eifler (2017), desilike [^2] https://github.com/cosmodesi/desilike, CosmoSIS Zuntz et al. (2015), Cobaya Torrado and Lewis (2021), firecrown [^3] https://github.com/LSSTDESC/firecrown, CLOE [^4] Euclid Collaboration in prep.). With each collaboration preferring different pipelines, performing joint analyses require substantial effort and computational resources – even when the analyses are statistically independent. It is this problem which we set out to solve.
@@
Explicitly sampling the likelihood requires the likelihood of each experiment to be integrated into a single unified likelihood package (e.g. CosmoSIS). This time consuming exercise may have to be repeated multiple times for each new experiment. The computational costs are also significant. A single DES Year 3 (DESY3) $3\times 2$ point analysis takes approximately 4 node days on a modern high-performance node with 128 cores Lemos et al. (2023).
Importance sampling offers a promising alternative. Suppose that we have access to samples from the posterior (chains) of one the experiments, $\{c_{i}\}$, over cosmological parameter space $\mathbb{c}$. Then to sample from the joint posterior we compute the likelihood of the second experiment at each point in the original chain $\{\mathcal{L}(c_{i})\}$ and then re-weight the original chain by these likelihoods. However, if both sets of experiments contain a large set of nuisance parameters, this option is no longer viable. This is because at every point in the first chain, there is a distribution of nuisance parameter values in the second experiment.
@@
In this paper we advocate for a new approach using normalizing flows. Given two chains from two independent cosmological experiments our approach is capable of drawing $\mathcal{O}(10^{9})$ from the joint posterior in seconds.
@@ II.2 Estimating the Marginal Posterior with Normalizing Flows
Given samples from a target distribution $p_{\mathbb{t}}(\mathbb{x})$, one can train a normalizing flow to learn an invertible mapping, $F$, from a multivariate normal distribution centered at the origin with identity covariance, to the target. As described in the previous section, we train normalizing flows to estimate the marginal posterior, $p(\mathbb{c}|\mathbb{d_{i}})$, from the chains.
@@ II.4 Using Outputs from Nested Samplers
Nested samplers such as PolyChord Handley et al. (2015), nautilus Lange (2023) and MultiNest Feroz et al. (2009) return a set of samples and corresponding weights. Since flowjax does not take weights as input, we use a custom loss function in CombineHarvesterFlow which uses the weighted average of the log-probabilities over the training set, rather than the average, used in flowjax
@@ IV.2 Validation of CombineHarvesterFlow and Causes of Failure
Even in the extreme examples presented in this section, CombineHarvesterFlow extracts reasonably accurate posteriors – while the results from the two different weighting strategies give a rough estimate on the uncertainties. In many circumstances (e.g. non-publishable testing and validation runs) this accuracy is more than sufficient, so we expect CombineHarester to dramatically accelerate workflows even when it fails to produce extremely accurate joint posteriors.
@@ IV.3 DESY3 $\times$ CMB Lensing + Boss Full Shape Direct Fit
Due to the different coverage in radial modes, the covariance between projected statistics and spectroscopic full shape statistics is negligible Taylor and Markovič (2022). Hence, we use CombineHarvesterFlow to derive joint constraints between the DESY3 $\times$ CMB lensing $6\times 2$ point posterior666publicly available at https://des.ncsa.illinois.edu/releases/y3a2 Abbott et al. (2023) and the BOSS full shape direct fit posteriors found in Philcox and Ivanov (2022) [^7] We use the chains which exclude the bispectrum information. To our knowledge this is the first time that this combination of probes has been combined into a single analysis. We pay particular attention to $S_{8}$ where tensions between CMB and large scale structure measurements have previously been reported (see e.g. Fig. 12 in Lange et al. (2023)).
We use CombineHarvesterFlow [^8] We trained an ensemble of 50 flows and checked that our results are converged. to find joint constraints in $\{\Omega_{\rm m},h_{0},\omega_{\rm b},n_{\rm s},S_{8}\}$. We take flat uninformative priors on $h_{0}$ and $A_{\rm s}$. Since the DESY3 $\times$ CMB lensing constraints ran up against the boundaries of the flat prior on $n_{\rm s}$, we impose a conservative Gaussian prior, $n_{\rm s}\sim\mathcal{N}(0.96,0.2)$, by retroactively weighting the chains after running CombineHarvesterFlow to avoid boundary effects. The DES chains were parameterised in $\{\Omega_{\rm m},h_{0},\Omega_{\rm b},n_{\rm s},A_{\rm s}\}$ space while the BOSS chains were run in $\{\omega_{\rm c},h_{0},\omega_{\rm b},n_{\rm s},\ln(10^{10}A_{\rm s})\}$. Hence, the contribution of the Jacobians cancel when going to $\{\Omega_{\rm m},h_{0},\omega_{\rm b},n_{\rm s},A_{\rm s}\}$-space, when we introduce a weighting factor of $A_{\rm s}$ to the chains. Both chains, store the $\sigma_{8}$ values at all sampled points, so we train the flows and present our final results in $S_{8}$, rather than $A_{\rm s}$. Finally, we take a Gaussian prior on $\omega_{\rm b}\sim N(0.02268,0.00036)$ as in the BOSS chains. Since the DES prior is uninformative, this is achieved simply by not dividing out by the $\omega_{\rm b}$ prior in Eqn. 2.
@@ VI Conclusion
CombineHarvesterFlow is fast. We have found that training time for each flow can vary depending on the hardware and size of parameters space – but training typically takes less than a minute. Once trained, the flows enable virtually instantaneous sampling of the joint posterior.
@@
CombineHarvesterFlow helps the user detect spurious results. In particular, weighting the first chain by the flows trained on the second chain, and vice versa, should yield consistent constraints.
@@ VII Acknowledgements
PLT is supported in part by NASA ROSES 21-ATP21-0050. This work received support from the U.S. Department of Energy under contract number DE-SC0011726. AC acknowledges support provided by NASA through the NASA Hubble Fellowship grant HST-HF2-51526.001-A awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Incorporated, under NASA contract NAS5-26555. The authors thank Mikhail Ivanov for providing the BOSS full shape chains used in this work. We acknowledge use of the open source software Bradbury et al. (2018); Hinton (2016); Lewis (2019); Virtanen et al. (2020); Harris et al. (2020); Hunter (2007)
@@ Appendix A Failure of Neural Network Likelihood Emulation
However, we find that direct emulation of the likelihood, rather than data-vector fails catastrophically. This is because most of the prior volume is not sampled by the chains leaving the emulated likelihood completely unconstrained in these regions. [^11] Emulating the data-vector does not suffer from this issue because in unconstrained regions the $\chi^{2}$ is most likely large and hence rejected by the sampler.
To illustrate this, we trained a likelihood emulator using the publicly available DESY3 $3\times 2$ point chains. Using this emulator we reran the DES chains, sampling with Emcee Foreman-Mackey et al. (2013). The results in the $\Omega_{\rm m}-A_{\rm s}$ plane, after marginalizing over the other parameters, are shown in Fig. 4. Here it is clear the emulated likelihood analysis fails catastrophically.

### v2

@@ Abstract
We show how to efficiently sample the joint posterior of two non-covariant experiments with a large set of nuisance parameters. Specifically, we train an ensemble of normalizing flows to learn the posterior distribution of both experiments. Once trained, we can use the flows to reweight $\mathcal{O}(10^{9})$ samples from both measurements to compute the joint posterior in seconds – saving up to $\mathcal{O}(1)$ ton of $\text{CO}_{2}$ per Monte Carlo run. Using this new technique we find joint constraints between the Dark Energy Survey $3\times 2$ point measurement, South Pole Telescope and Planck CMB lensing and a BOSS direct fit full shape analyses, for the first time. We find $\Omega_{\rm m}=0.32^{+0.01}_{-0.01}$ and $S_{8}=0.79^{+0.01}_{-0.01}$. We release a public package called CombineHarvesterFlow [^1] https://github.com/pltaylor16/CombineHarvesterFlow which performs these calculations.
@@ I Introduction
Analyses of data from the Atacama Cosmology Telescope (ACT) Qu et al. (2024), the Baryon Oscillation Spectroscopic Survey (BOSS) Alam et al. (2021), the Dark Energy Survey (DES) Abbott et al. (2022), the Dark Energy Spectroscopic Instrument (DESI) Adame et al. (2024), the extended Baryon Oscillation Spectroscopic Survey (eBOSS) Alam et al. (2021), Hyper Suprime-Cam (HSC) More et al. (2023), the Kilo-degree Survey (KiDS) Heymans et al. (2021), Planck Aghanim et al. (2020) and the South Pole Telescope (SPT) Bianchini et al. (2020), have provided some of the tightest cosmological constraints. Future measurements with Euclid Mellier et al. (2024), the Nancy Grace Roman Space Telescope [^2] https://roman.gsfc.nasa.gov/, the Rubin Observatory [^3] https://www.lsst.org/, the Simons Observatory [^4] https://simonsobservatory.org/ and CMB-S4555https://cmb-s4.org/ will tighten these constraints even further.
Meanwhile there has been a proliferation of cosmological likelihood packages to perform inference with these datasets (e.g. MontePython Brinckmann and Lesgourgues (2019), CosmoMC Lewis (2013), COSMOLIKE Krause and Eifler (2017), desilike [^6] https://github.com/cosmodesi/desilike, CosmoSIS Zuntz et al. (2015), Cobaya Torrado and Lewis (2021), firecrown [^7] https://github.com/LSSTDESC/firecrown, CLOE [^8] Euclid Collaboration in prep.). With each collaboration preferring different pipelines, performing joint analyses requires substantial effort and computational resources – even when the analyses are statistically independent. It is this problem which we set out to solve.
@@
Explicitly sampling the likelihood requires the likelihood of each experiment to be integrated into a single unified likelihood package (e.g. CosmoSIS). This time consuming exercise may have to be repeated multiple times for each new experiment. The computational costs are also significant. A single DES Year 3 (DESY3) $3\times 2$ point analysis takes up to 4 node days on a modern high-performance node with 128 cores Lemos et al. (2023).
Importance sampling offers a promising alternative. Suppose that we have access to samples from the posterior (chains) of one of the experiments, $\{c_{i}\}$, over cosmological parameter space $\mathbb{c}$. Then to sample from the joint posterior we compute the likelihood of the second experiment at each point in the original chain $\{\mathcal{L}(c_{i})\}$ and then re-weight the original chain by these likelihoods. However, if both sets of experiments contain a large set of nuisance parameters, this option is no longer viable. This is because at every point in the first chain, there is a distribution of nuisance parameter values in the second experiment.
@@
In this paper we advocate for a new approach using normalizing flows. Given two chains from two independent cosmological experiments our approach is capable of reweighting $\mathcal{O}(10^{9})$ points from each experiment to compute the joint posterior in seconds.
@@ II.2 Estimating the Marginal Posterior with Normalizing Flows
Given samples from a target distribution $p_{\mathbb{t}}(\mathbb{x})$, one can train a normalizing flow to learn an invertible mapping, $F$, from a multivariate normal distribution centered at the origin with identity covariance, to the target. Once the mapping is learned, one can estimate the likelihood of each point in the target from the Jacobian: $p_{\mathbb{t}}(\mathbb{x})=\mathcal{G}(F^{-1}(\mathbb{x}))|\frac{\partial F^{-1}}{\partial x}|$, where $\mathcal{G}$ is the probability density of the unit normal. We use this feature of normalizing flows to estimate the marginal posterior, $p(\mathbb{c}|\mathbb{d_{i}})$, from the chains as described in the previous section.
@@ II.4 Using Outputs from Nested Samplers
Nested samplers such as PolyChord Handley et al. (2015), nautilus Lange (2023) and MultiNest Feroz et al. (2009) return a set of samples and corresponding weights. Since flowjax does not take weights as input, we use a custom loss function in CombineHarvesterFlow which uses the weighted average of the log-probabilities over the training set, rather than the average, used in flowjax.
@@ IV.2 Validation of CombineHarvesterFlow and Causes of Failure
Even in the extreme examples presented in this section, CombineHarvesterFlow extracts reasonably accurate posteriors – while the results from the two different weighting strategies give a rough estimate on the uncertainties. In many circumstances (e.g. non-publishable testing and validation runs) this accuracy is more than sufficient, so we expect CombineHaresterFlow to dramatically accelerate workflows even when it fails to produce extremely accurate joint posteriors. We stress that these errors are typically much smaller than those expected from model mis-specification. In the future, we recommend that users specify a success criterion (e.g. fractional error on the error for parameters of interest) and add the differences from the CombineHaresterFlow cross-check to the total error budget.
@@ IV.3 DESY3 $\times$ CMB Lensing + Boss Full Shape Direct Fit
Due to the different coverage in radial modes, the covariance between projected statistics and spectroscopic full shape statistics is negligible Taylor and Markovič (2022). Hence, we use CombineHarvesterFlow to derive joint constraints between the DESY3 $\times$ CMB lensing $6\times 2$ point posterior101010publicly available at https://des.ncsa.illinois.edu/releases/y3a2 Abbott et al. (2023) and the BOSS full shape direct fit posteriors found in Philcox and Ivanov (2022) [^11] We use the chains which exclude the bispectrum information. To our knowledge this is the first time that this combination of probes has been combined into a single analysis. We pay particular attention to $S_{8}$ where tensions between CMB and large scale structure measurements have previously been reported (see e.g. Fig. 12 in Lange et al. (2023)).
We use CombineHarvesterFlow [^12] We trained an ensemble of 50 flows and checked that our results are converged. to find joint constraints in $\{\Omega_{\rm m},h_{0},\omega_{\rm b},n_{\rm s},S_{8}\}$. We take flat uninformative priors on $h_{0}$ and $A_{\rm s}$. [^13] We note however, that our choice of $n_{s}$ prior can have a substantial impact on the inferred value of $h_{0}$ in CMB lensing Kable et al. (2023). Since the DESY3 $\times$ CMB lensing constraints ran up against the boundaries of the flat prior on $n_{\rm s}$, we impose a conservative Gaussian prior, $n_{\rm s}\sim\mathcal{N}(0.96,0.02)$, by retroactively weighting the chains after running CombineHarvesterFlow to avoid boundary effects. The DES chains were parameterised in $\{\Omega_{\rm m},h_{0},\Omega_{\rm b},n_{\rm s},A_{\rm s}\}$ space while the BOSS chains were run in $\{\omega_{\rm c},h_{0},\omega_{\rm b},n_{\rm s},\ln(10^{10}A_{\rm s})\}$. Hence, the contribution of the Jacobians cancel when going to $\{\Omega_{\rm m},h_{0},\omega_{\rm b},n_{\rm s},A_{\rm s}\}$-space, when we introduce a weighting factor of $A_{\rm s}$ to the chains. Both chains, store the $\sigma_{8}$ values at all sampled points, so we train the flows and present our final results in $S_{8}$, rather than $A_{\rm s}$. Finally, we take a Gaussian prior on $\omega_{\rm b}\sim N(0.02268,0.00036)$ as in the BOSS chains. Since the DES prior is uninformative, this is achieved simply by not dividing out by the $\omega_{\rm b}$ prior in Eqn. 2.
@@ VI Conclusion
CombineHarvesterFlow is fast. We have found that training time for each flow can vary depending on the hardware and size of parameters space – but training typically takes less than a minute on a CPU. Once trained, the flows enable virtually instantaneous sampling of the joint posterior.
@@
CombineHarvesterFlow helps the user detect spurious results. In particular, weighting the first chain by the flows trained on the second chain, and vice versa, should yield consistent constraints. After submission, we became aware of several papers which proposed similar techniques to the one proposed in this work Bevins et al. (2022, 2023). Subsequent to this work Mootoovaloo et al. (2024) also proposed a similar method. However, the ability to detect spurious results with the cross-check is a key feature that sets CombineHarvesterFlow apart.
@@ VII Acknowledgements
PLT is supported in part by NASA ROSES 21-ATP21-0050. This work received support from the U.S. Department of Energy under contract number DE-SC0011726. AC acknowledges support provided by NASA through the NASA Hubble Fellowship grant HST-HF2-51526.001-A awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Incorporated, under NASA contract NAS5-26555. The authors thank Mikhail Ivanov for providing the BOSS full shape chains used in this work and the anonymous referee who’s comments improved the manuscript. We acknowledge use of the open source software Bradbury et al. (2018); Hinton (2016); Lewis (2019); Virtanen et al. (2020); Harris et al. (2020); Hunter (2007)
@@ Appendix A Failure of Neural Network Likelihood Emulation
However, we find that direct emulation of the likelihood, rather than data-vector can fail catastrophically. This is because most of the prior volume is not sampled by the chains leaving the emulated likelihood completely unconstrained in these regions. [^16] Emulating the data-vector does not suffer from this issue because in unconstrained regions the $\chi^{2}$ is most likely large and hence rejected by the sampler.
To illustrate this, we trained a likelihood emulator using the publicly available DESY3 $3\times 2$ point chains. Using this emulator we reran the DES chains, sampling with Emcee Foreman-Mackey et al. (2013). The results in the $\Omega_{\rm m}-A_{\rm s}$ plane, after marginalizing over the other parameters, are shown in Fig. 4. Here it is clear the emulated likelihood analysis has failed. It may be possible to solve this problem by (e.g. forcing the posterior to zero away from the samples), but this could require substantial fine-tuning to the problem at hand.